Grade 11Magical Mathematics[Interesting Approach]

The numbers have been arranged under some rule. Based on that rule which number will come on the place of Question Mark ?
1,1,2,6,24,?,720

Shri Krishan Baghel
10 Years agoGrade 11

3 Answers

Karthik Kumar
10 Years ago
1=1
1=1*1(multiply preceding number with 1)
2=1*2(multiply preceding number with 2)
6=2*3(multiply preceding number with 3)
24=6*4(multiply preceding number with 4)
?=24*5=(120)(multiply preceding number with 5 .then  answer is  120)
720=120*6 (cleared answer is 120
multiply preceding number with 1.2.3.4.5.6  order
NARAYANA
10 Years ago
MY DEAR FRIEND YOU ASKED LOGICAL THINKING QUASTION DO THESE TYPE OF QUASTION SOME TIME IS REQUIRED AT LAST  IT’S ANSWER IS 120.THANK YOU.
KALYAN
10 Years ago
answer is 120 because 24*5 =  120
the logic is 1*1
1*2
2*3 in these way the series goes on 6*4 = 24 then 24*5 = 120
so answer is 120